# Dynamics-Aware Gaussian Splatting Streaming Towards Fast On-the-Fly Training for 4D Reconstruction

### v1

@@ 4.1 Experimental Setup
Baselines. Our primary comparisons are against iterative 4D dynamic spatial reconstruction methods that support on-the-fly training and per-timestep streaming, specifically StreamRF [25] and 3DGStream [38], which are categorized as online methods. To evaluate reconstruction fidelity, we also include other 4D dynamic spatial reconstruction methods that require full-length multi-view videos as input and do not support on-the-fly training or per-timestep streaming. This offline category includes NeRF-based methods such as DyNeRF [26], NeRFPlayer [37], HexPlane [5], K-Planes [14], HyperReel [2], and MixVoxels [42], as well as recent 3DGS-based methods like 4DGS-Wu [47], 4DGS-Yang [52], E-D3DGS [3], and STG [27]. Note that NeRFPlayer [37] allows streaming after optimization but does not support on-the-fly training. In addition, we consider static methods, including Plenoxels [12] and I-NGP [34], which fully train a static scene representation for each frame.
@@
Metrics. We focus on achieving fast on-the-fly training for streamable dynamic spatial reconstruction, emphasizing both time efficiency and reconstruction fidelity. To evaluate time efficiency, we calculate the average per-frame training time over the 300 frames in the video and measure the frames per second (FPS). For offline methods that do not support on-the-fly training, we report their training time averaged over all frames. To assess reconstruction fidelity, we use metrics including peak signal-to-noise ratio (PSNR) and dissimilarity structural similarity index measure (DSSIM).
@@
Table 2: Quantitative comparison on the Meet Room dataset. The training time and quality metrics are averaged over all 300 frames.
@@
StreamRF [25] 26.72 10.2 10
3DGStream [38] 30.79 6 288
DASS (Ours) 30.81 5.04 253.8
Figure 5: Qualitative comparison for the flame steak scene of the N3DV dataset, including online training methods [25, 38] and offline reconstruction method [47]. Compared to other methods, our method accurately recovers dynamic objects and preserves intricate details.
4.2 Results and Comparisons
The quantitative evaluation results of our method on the N3DV dataset [26] are detailed in Tab. 1, and a comprehensive comparison of multiple metrics is shown in Fig. 1. Our method achieves the fastest on-the-fly training speed, which converges from previous frame to the current one within 10 seconds, yielding a 20% improvement in efficiency. In contrast, the average training time of baseline methods ranges from tens of seconds to several minutes. This time efficiency advantage originates from several aspects: First, our inheritance stage effectively captures the temporal continuity and selectively preserves the optimized results from the previous frame. This significantly alleviates the optimization burden in the following stages, especially for the densification stage, which requires 60 steps compared to the 100 steps in 3DGStream. Moreover, the selective inheritance stage is lightweight, as it optimizes only one learnable parameter for each Gaussian primitive, in contrast to the tens of parameters in a pipeline fully dependent on densification. In addition, our dynamics-aware shift stage assigns different deformation layers for dynamic and static components, which focus on significant and subtle variations, respectively. This strategy enhances the convergence achieving high-quality deformations with fewer optimization steps (100 steps instead of 150 steps in 3DGStream). Although our pipeline requires optical flow and segmentation, these results are either directly inferred or rendered within milliseconds and are executed only once every few frames.
Beyond the significant time advantage, the reconstruction fidelity of our method also outperforms previous online streamable baselines [25, 38] and surpasses most state-of-the-art offline training methods [26, 37, 5, 14, 2, 42, 47, 3]. Our dynamics-aware deformation places greater emphasis on areas with complex motions and variations, thereby facilitating more detailed reconstruction. Besides, our error-guided densification strategy identifies areas with weak construction and newly emerging objects, which rapidly detects and mitigates the reconstruction defects. A qualitative comparison on the N3DV dataset is provided in Fig. 5. A detailed comparison of each scene in the N3DV dataset is provided in the supplementary material.
@@ 4.2 Results and Comparisons
Method PSNR(dB)$\uparrow$
w/o Inheritance 31.71
w/o Dynamics awareness 31.69
w/o Densification 31.50
w/o Error guidance 31.88
Full Model (Our DASS) 31.99
Our rendering speed further highlights the superiority of our method over others. Leveraging the real-time rasterization, our rendering speed significantly outperforms that of NeRF-based methods. Moreover, our method achieves higher FPS than 4DGS methods, since our deformation fields, comprising only hash encoding and simple MLPs, enable a low query overhead of around 1 ms.
4.3 Ablation Studies
We conduct ablation experiments on the N3DV dataset to evaluate the contributions of our proposed designs. Our pipeline comprises three key stages: selective inheritance, dynamics-aware shift, and error-guided densification. To assess the impact of each stage, we either replace it with a baseline counterpart or remove it from the pipeline, and provide reconstruction quality comparisons with fixed optimization steps. The results are summarized in Tab. 3.
Effectiveness of Selective Inheritance. To evaluate the impact of the selective inheritance stage, we remove it from the pipeline, leading to the baseline “w/o Inheritance” in Tab. 3. In this baseline, base Gaussians are directly transformed, and newly added Gaussian are regenerated for each frame. This approach results in a performance degradation of 0.28 dB compared to the full model, since it fails to leverage the temporal continuity and imposes optimization burdens on the following two stages. In contrast, our selective inheritance stage effectively balances the preservation of essential Gaussians while controlling numerical accumulation. We also visualize the error maps before and after the inheritance stage in Fig. 6, which demonstrates that our selective inheritance strategy effectively enhances temporal coherence and reduces the optimization overhead for subsequent stages.
Effectiveness of Dynamics-Aware Shift. To validate the improvement of our dynamics-aware shift strategy, we replace the dynamics-aware deformation fields with one universal hash-encoding MLP, referred to as “w/o Dynamics awareness” in Tab. 3. This baseline shows a performance drop of 0.3 dB, which underscores the importance of our dynamics-aware optimization in highlighting the intrinsic dynamics and statics in the space and effectively managing the complex movements.
Figure 6: Error map visualization without (left) and with (right) the proposed selective inheritance mechanism. The selective inheritance stage effectively compensates for the artifacts in temporally consistent area and mitigates the optimization distortions.
Effectiveness of Error Guided Densification. To evaluate the effectiveness of our error-guided densification, we consider two baselines for comparison. The first “w/o Densification” baseline excludes the densification stage and results in a significant decline in quality. This suggests that solely relying on transformations of existing Gaussian primitives is inadequate for representing dynamic scenes. The second baseline, “w/o Error guidance”, removes the error-guided subset $\mathcal{S}_{\textrm{err}}$ from the densification process and uses only the gradient indicator, leading to degraded performance. This indicates that our error-guided densification effectively caters for under-reconstructed areas and emerging objects.
@@ 5 Conclusion
In this paper, we introduce a novel three-stage pipeline for iterative online 4D dynamic spatial reconstruction, which allows for on-the-fly training and per-frame streaming. We incorporate a selective inheritance stage to capture the temporal continuity, a dynamics-aware shift stage to exploit the dynamic and static features in natural scenes, and an error-guided densification stage to adaptively recover emerging objects and weak reconstruction. Our method achieves state-of-the-art performance in online streaming 4D reconstruction, providing the fastest training speed, with a 20% improvement, and superior representation quality.
@@
The supplementary material provides an introduction of preliminary knowledge in Sec. A, more implementation details in Sec. B, additional experimental results in Sec. C, and discusses potential extensions in Sec. D.
@@ B.1 Implementation and Metrics
We implement our streaming pipeline based on the official open-source codebase 3DGStream and strictly follow the evaluation principles. Specifically, the reported training time includes the time consumption for training the initialization at frame 0. This frame 0 initialization process takes 10.18 minutes in our pipeline and is counted into the average training time calculation following the official 3DGStream. Notably, the time consumption of subsequent on-the-fly training takes only 7.59 s for each frame. The experiments are mainly conducted on an RTX A6000 GPU and an Intel Xeon Platinum 8370C CPU.
@@ B.2 Additional Details in Each Stage
Initialization. For the initialization at frame 0, we employ the official codebase of Gaussian Grouping. Following 3DGStream, the spherical harmonics degree is set to 1. The segmentation-related hyperparameters in Gaussian Grouping remain unmodified, as we find that varying these hyperparameters leads to minimal impact on the streaming performance. Moreover, during training the initialization, we introduce an opacity regularization term, written as $-\frac{1}{N}\sum^{N}_{i=1}o_{i}\log(o_{i})$, in addition to the fidelity loss. This opacity regularization encourages the opacity of Gaussian primitives to approach either zero or one, thereby pushing Gaussians to the object surface and naturally pruning nearly transparent Gaussians. This regularization has proven effective in previous works [10, 54] to enhance the representation efficiency and quality.
Selective Inheritance. In our proposed selective inheritance stage, the quantized learnable vector after the sigmoid function, $\textrm{Quant}(\textrm{sigmoid}(\mathbf{m}))$, determines whether to maintain the densified Gaussian primitives from the previous frame. Notably, the quantization operation itself is non-differentiable. Therefore, the Straight Through Estimator (STE) [4] is applied in the optimization of $\mathbf{m}$, which is written as:
@@
where $\mathbf{m}_{\textrm{op}}$ is the parameter involved in differentiable optimization. Moreover, this selective inheritance is paused every 20 frames, i.e., the inherited Gaussians are erased once for every 20 frames. This strategy considers the typical duration of temporally consistent objects and avoids introducing misleading information during optimization.
Dynamics-Aware Shift. The dynamics-aware shift stage distinguishes the dynamic and static groups of Gaussian primitives in the representation and employs different hash-encoding MLPs, $\mathcal{H}_{\textrm{dyn}}$ and $\mathcal{H}_{\textrm{st}}$, as deformation fields for their optimization. The estimation of dynamic objects usually remains consistent across frames, so repeatedly performing this estimation introduces temporal redundancy into the on-the-fly training. In our pipeline, the estimation of dynamic Gaussian group (as shown in Fig. 4 in the main body of the paper) is performed every 10 frames, considering that the dynamic properties remain stable across adjacent frames. As discussed in Sec. 3.3, the hash table size $T_{\textrm{Hash}}$ and the number of feature dimensions per entry $F_{\textrm{Hash}}$ vary between dynamic and static groups. Specifically, for the N3DV dataset, the dynamic group adopts $T_{\textrm{Hash}}=2^{16}$ and $F_{\textrm{Hash}}=4$ for $\mathcal{H}_{\textrm{dyn}}$, while the static group uses $T_{\textrm{Hash}}=2^{14}$ and $F_{\textrm{Hash}}=2$ for $\mathcal{H}_{\textrm{st}}$. For the Meet Room dataset, the dynamic group employs $T_{\textrm{Hash}}=2^{15}$ and $F_{\textrm{Hash}}=4$, while the static group adopts $T_{\textrm{Hash}}=2^{13}$ and $F_{\textrm{Hash}}=2$.
@@ C.1 Additional Ablation Studies
Impact of Optimization Steps. To assess the impact of the optimization steps at each stage and illustrate the convergence efficiency of our DASS framework, we adjust the optimization steps and evaluate the reconstruction qualities, as shown in Tab. 4. Our DASS assigns $s_{1}=20$, $s_{2}=100$, and $s_{3}=60$ for the inheritance, shift, and densification stages, respectively, serving as the baseline in the first row of Tab. 4. This assignment contributes to the time efficiency of our method compared to the state-of-the-art baseline 3DGStream [38] (9.62 s compared to 12 s), which assigns a total of 250 optimization steps (150 for transformation and 100 for densification). The baselines with increased optimization steps shown in subsequent rows verify that our DASS achieves optimal convergence with the primary optimization setting. Increasing the optimization steps leads to a negligible improvement in representation quality, less than 0.05 dB, while significantly increasing time consumption, thereby validating our convergence efficiency. This efficiency comes from multiple aspects: Our selective inheritance stage preserves the essential Gaussian primitives of the previous frame’s densified Gaussians and mitigates the optimization distortions with only $s_{3}=20$ steps of optimization. The dynamics-aware shift stage learns the diverse distributions of dynamic and static components, respectively, facilitating the deformation convergence. Besides, assigning a low-complexity deformation field to the static components conserves temporal and computational resources for the majority of the Gaussian primitives. The error-guided densification stage effectively detects and compensates for the distortions, enhancing the reconstruction quality while introducing low temporal overhead. In addition, the last two baselines in Tab. 4 demonstrate that reducing the training steps leads to inferior reconstruction quality, necessitating our optimization step assignment.
@@ C.2 Quantitative Results
In addition to the state-of-the-art 4D reconstruction baselines shown in Tab. 1 in the main body of the paper, we provide comparisons with two variants of naive 3DGS methods in Tab. 5 and Fig. 7. Specifically, we compare our method with a naive streamable baseline, which trains an independent set of 3D Gaussian primitives for each frame, referred to as “Naive Per-Frame 3DGS” in Tab. 5. In this baseline, each frame is trained with the same conditions as the initial zero-frame in streaming methods. Although this naive baseline outperforms all off-the-shelf baselines in 4D reconstruction, it fully trains a set of Gaussian primitives from scratch for each scene, leading to a significant training time overhead of approximately 615 seconds per-frame. In contrast, our DASS effectively models the temporal variations in the 3D space and facilitates a fast convergence between frames. DASS demonstrates competitive streaming quality, with only a negligible gap of 0.08 dB compared to this naive baseline, while consuming merely 1.56% of the training time. This significant improvement highlights the efficiency of DASS. We also compare our method with an intuitive full-tuned baseline, which iteratively tunes all parameters of the Gaussian primitives initialized from the previous timestep, termed “Intuitive Full-Tuned 3DGS” in Tab. 5. Specifically, it initializes the on-the-fly training with the same frame 0 initialization as our DASS and tunes all Gaussian properties for a fixed number of iterations, which yields a similar time consumption to ours. As shown in Tab. 5, a significant performance degradation of 1.16 dB is observed in this baseline. This indicates that the intuitive baseline fails to capture temporal variations between frames under such time efficiency, while our DASS effectively manages and converges these variations with lower temporal overhead.

### v2

@@ 4.1 Experimental Setup
Baselines. Our primary comparisons are against iterative 4D dynamic spatial reconstruction methods that support on-the-fly training and per-timestep streaming, specifically StreamRF [25], 3DGStream [37] and HiCoM [15], which are categorized as online methods. To ensure fair comparisons, we evaluate these online methods on the same RTX A6000 GPU. To evaluate reconstruction fidelity, we also include other 4D dynamic spatial reconstruction methods that require full-length multi-view videos as input and do not support on-the-fly training or per-timestep streaming. This offline category includes NeRF-based methods such as DyNeRF [26], NeRFPlayer [36], HexPlane [5], K-Planes [14], HyperReel [2], and MixVoxels [41], as well as recent 3DGS-based methods like 4DGS-Wu [46], 4DGS-Yang [51], E-D3DGS [3], and STG [27]. Note that NeRFPlayer [36] allows streaming after optimization but does not support on-the-fly training. In addition, we consider static methods, including Plenoxels [12] and I-NGP [33], which fully train a static scene representation for each timestep.
@@
Metrics. We focus on fast on-the-fly training for streamable dynamic spatial reconstruction, emphasizing both time efficiency and reconstruction fidelity. To evaluate time efficiency, we calculate the average per-timestep training time across the entire video sequence and measure the frames per second (FPS). For offline methods that do not support on-the-fly training, we report their training time averaged over all timesteps. To assess reconstruction fidelity, we use metrics including peak signal-to-noise ratio (PSNR) and dissimilarity structural similarity index measure (DSSIM).
4.2 Results and Comparisons
The quantitative evaluation results of our method on the N3DV dataset [26] are detailed in Tab. 1, and a comprehensive comparison of multiple metrics is shown in Fig. 1 (Middle). Our method achieves the fastest on-the-fly reconstruction speed, which converges from the previous timestep to the current one within 10 seconds, yielding a significant improvement in efficiency. In contrast, the average training time of baseline methods ranges from tens of seconds to several minutes. This significant time efficiency advantage originates from several key designs: First, our inheritance stage effectively captures the temporal continuity and selectively preserves the optimized results from the previous timestep. This significantly alleviates the optimization burden in subsequent stages, especially in the densification stage, which requires only 60 steps compared to the 100 steps in 3DGStream. Moreover, the selective inheritance stage is highly lightweight, as it optimizes only one learnable parameter for each Gaussian primitive, in contrast to the tens of parameters required in a fully tuned pipeline. In addition, our dynamics-aware shift stage assigns different deformation fields for dynamic and static components, focusing on significant and subtle variations, respectively. This strategy enhances the convergence and achieves high-quality deformations with fewer optimization steps (100 steps instead of 150 steps in 3DGStream). Although our pipeline requires optical flow and segmentation, these results are either directly inferred or rendered within milliseconds and are executed only once every few timesteps.
Beyond the significant efficiency advantage, our reconstruction fidelity also outperforms previous online streamable baselines [25, 37, 15] and surpasses most state-of-the-art offline training methods [26, 36, 5, 14, 2, 41, 46, 3]. Our dynamics-aware deformation places greater emphasis on areas with complex motions and variations, thereby facilitating more detailed reconstruction. Besides, our error-guided densification strategy identifies areas with weak construction and newly emerging objects, which rapidly detects and mitigates the reconstruction defects. Qualitative comparisons on the N3DV dataset, along with a visualization of our detected dynamics components, are provided in Fig. 6 (Left). A detailed comparison of each scene on the N3DV dataset is provided in the supplementary material.
Our rendering speed further highlights the superiority of our method over others. Leveraging the real-time 3DGS rasterization, our rendering speed significantly outperforms that of NeRF-based methods. Moreover, our method achieves higher FPS than 4DGS methods, since our deformation fields, comprising only hash encoding and simple MLPs, enable a low query overhead of around 1 ms.
@@ 4.2 Results and Comparisons
Table 2: Quantitative comparison on the Meet Room dataset, with metrics averaged over all frames. Following previous baselines, the scene Discussion is chosen for evaluation.
@@
StreamRF [25] 27.03 9.97 12
3DGStream [37] 30.79 6.20 275
HiCoM [15] 29.87 5.58 277
DASS (Ours) 31.02 4.91 254
4.3 Ablation Studies
We conduct ablation experiments on the N3DV dataset to evaluate the contributions of each component in our proposed design. We start from a baseline optimizing Gaussian positions and progressively transition to our full model, with all variants maintaining consistent optimization steps. The results regarding training time and reconstruction fidelity are summarized in Tab. 3, which demonstrates the effectiveness of our designs in improving reconstruction quality while minimizing training time. Additional ablation studies are provided in Supplementary Sec. 8.1.
Effectiveness of Dynamics-Aware Shift. To validate the improvement brought by our dynamics-aware shift strategy, we compare two variants: one with a uniform shift deformation field (+ Shift) and another with dynamics-aware deformation (+ Dynamics aware), as referenced in Tab. 3. The latter achieves a better PSNR performance with comparable time consumption, which underscores the importance of our dynamics-aware optimization in distinguishing between intrinsic dynamic and static components and effectively managing the complex movements.
Effectiveness of Error-Guided Densification. To evaluate the effectiveness of our error-guided densification, we consider three variants in Tab. 3. A plain densification stage, similar to pervious works [21, 37] (+ Densification), yields modest performance improvements. Simply lowering the densification threshold $\tau_{\textrm{pos}}$ (+ Lower threshold) results in an excessive number of Gaussians, increasing training time by 0.79s. In contrast, our proposed error-guided densification strategy (+ Error guidance), which integrates $\mathcal{S}_{\textrm{err}}$ as in Eq. 4, adaptively focuses on regions with weak reconstruction and emerging objects, providing superior performance.
Effectiveness of Selective Inheritance. To evaluate the impact of the selective inheritance stage, we compare two variants in Tab. 3. Directly inheriting all Gaussians from previous timestep (+ Direct inheritance) leads to excessive accumulation of Gaussians, which hinders fast on-the-fly optimization and introduces errors to the reconstruction. Conversely, our selective inheritance stage (+ Selective inheritance) effectively exploits the temporal consistency, balancing the preservation of essential Gaussians while controlling numerical accumulation. We also visualize the error maps before and after the selective inheritance stage in Fig. 6 (Right), which demonstrates that our selective inheritance strategy effectively recovers temporal coherence with highly lightweight optimization. By optimizing just one parameter per Gaussian for $s_{3}=20$ steps, our approach reduces the optimization overhead for subsequent stages.
@@ 4.3 Ablation Studies
Method PSNR (dB)$\uparrow$ Time (s)$\downarrow$ Method PSNR (dB)$\uparrow$ Time (s)$\downarrow$
Baseline 30.60 11.35 + Lower threshold 31.68 10.78
+ Shift 30.95 9.97 + Error guidance 31.76 10.03
+ Dynamics aware 31.31 10.08 + Direct inheritance 31.71 11.88
+ Densification 31.60 9.99 + Selective inheritance 31.99 9.62
@@ 5 Conclusion
In this paper, we introduce a novel three-stage pipeline for iterative online 4D dynamic spatial reconstruction, which allows for on-the-fly training and per-timestep streaming. We incorporate a selective inheritance stage to capture the temporal continuity, a dynamics-aware shift stage to exploit the dynamic and static features in natural scenes, and an error-guided densification stage to adaptively recover emerging objects and weak reconstruction. Our method achieves state-of-the-art performance in online streaming 4D reconstruction, providing the fastest training speed and superior representation quality.
@@
The supplementary material provides an introduction of preliminary knowledge in Sec. A, more implementation details in Sec. B, additional experimental results in Sec. C, analysis on concurrent works in Sec. D, and discusses potential extensions in Sec. E.
@@ B.1 Implementation and Metrics
We implement our streaming pipeline based on the official open-source codebase of 3DGStream [37] and strictly follow the evaluation principles. Specifically, the reported training time includes the time consumption for training the initialization at frame 0. This timestep 0 initialization process takes 10.18 minutes in our pipeline and is counted into the average training time calculation following the official setup of 3DGStream. Notably, the time consumption of subsequent on-the-fly training takes only 7.59 s for each timestep. The experiments are conducted on an RTX A6000 GPU and an Intel Xeon Platinum 8370C CPU.
@@ B.2 Additional Details in Each Stage
Initialization. For the initialization at timestep 0, we employ the official codebase of Gaussian Grouping. The segmentation-related hyperparameters in Gaussian Grouping remain unmodified, as we find that varying these hyperparameters leads to minimal impact on the streaming performance. Moreover, during training the initialization, we introduce an opacity regularization term, written as $-\frac{1}{N}\sum^{N}_{i=1}o_{i}\log(o_{i})$, in addition to the fidelity loss. This opacity regularization encourages the opacity of Gaussian primitives to approach either zero or one, thereby pushing Gaussians to the object surface and naturally pruning nearly transparent Gaussians. This regularization has proven effective in previous works [9, 54] to enhance the representation efficiency and quality.
Selective Inheritance. In our proposed selective inheritance stage, the quantized learnable vector after the sigmoid function, $\textrm{Quant}(\textrm{sigmoid}(\mathbf{m}))$, determines whether to maintain the densified Gaussian primitives from the previous timestep. Notably, the quantization operation itself is non-differentiable. Therefore, the Straight Through Estimator (STE) [4] is applied in the optimization of $\mathbf{m}$, which is written as:
@@
where $\mathbf{m}_{\textrm{op}}$ is the parameter involved in differentiable optimization. Moreover, this selective inheritance is paused every 20 timesteps, i.e., the inherited Gaussians are erased once for every 20 timesteps. This strategy considers the typical duration of temporally consistent objects and avoids introducing misleading information during optimization.
Dynamics-Aware Shift. The dynamics-aware shift stage distinguishes the dynamic and static groups of Gaussian primitives in the representation and employs different hash-encoding MLPs, $\mathcal{H}_{\textrm{dyn}}$ and $\mathcal{H}_{\textrm{st}}$, as deformation fields for their optimization. The estimation of dynamic objects usually remains consistent across timesteps, so repeatedly performing this estimation introduces temporal redundancy into the on-the-fly training. In our pipeline, the estimation of dynamic Gaussian group (as shown in Fig. 4 in the main body of the paper) is performed every 10 timesteps, considering that the dynamic properties remain stable across adjacent timesteps. As discussed in Sec. 3.3, the expressive ability of I-NGP is affected by the network complexity parameters, including the hash table size $T_{\textrm{Hash}}$ and the number of feature dimensions per entry $F_{\textrm{Hash}}$. Subsequently, the hash table size $T_{\textrm{Hash}}$ and the number of feature dimensions per entry $F_{\textrm{Hash}}$ vary between dynamic and static groups. Specifically, for the N3DV dataset, the dynamic group adopts $T_{\textrm{Hash}}=2^{16}$ and $F_{\textrm{Hash}}=4$ for $\mathcal{H}_{\textrm{dyn}}$, while the static group uses $T_{\textrm{Hash}}=2^{14}$ and $F_{\textrm{Hash}}=2$ for $\mathcal{H}_{\textrm{st}}$. For the Meet Room dataset, the dynamic group employs $T_{\textrm{Hash}}=2^{15}$ and $F_{\textrm{Hash}}=4$, while the static group adopts $T_{\textrm{Hash}}=2^{13}$ and $F_{\textrm{Hash}}=2$.
@@ C.1 Additional Ablation Studies
Impact of Optimization Steps. To assess the impact of the optimization steps at each stage and illustrate the convergence efficiency of our DASS framework, we adjust the optimization steps and evaluate the reconstruction qualities, as shown in Tab. 4. Our DASS assigns $s_{1}=20$, $s_{2}=100$, and $s_{3}=60$ for the inheritance, shift, and densification stages, respectively, serving as the baseline in the first row of Tab. 4. This assignment contributes to the time efficiency of our method compared to the state-of-the-art baseline 3DGStream [37] (9.62 s compared to 12 s), which assigns a total of 250 optimization steps (150 for transformation and 100 for densification). The baselines with increased optimization steps shown in subsequent rows verify that our DASS achieves optimal convergence with the primary optimization setting. Increasing the optimization steps leads to a negligible improvement in representation quality, less than 0.05 dB, while significantly increasing time consumption, thereby validating our convergence efficiency. This efficiency comes from multiple aspects: Our selective inheritance stage preserves the essential Gaussian primitives of the previous timestep’s densified Gaussians and mitigates the optimization distortions with only $s_{3}=20$ steps of optimization. The dynamics-aware shift stage learns the diverse distributions of dynamic and static components, respectively, facilitating the deformation convergence. Besides, assigning a low-complexity deformation field to the static components conserves temporal and computational resources for the majority of the Gaussian primitives. The error-guided densification stage effectively detects and compensates for the distortions, enhancing the reconstruction quality while introducing low temporal overhead. In addition, the last two baselines in Tab. 4 demonstrate that reducing the training steps leads to inferior reconstruction quality, necessitating our optimization step assignment.
@@ C.2 Quantitative Results
In addition to the state-of-the-art 4D reconstruction baselines shown in Tab. 1 in the main body of the paper, we provide comparisons with two variants of naive 3DGS methods in Tab. 5 and Fig. 7. Specifically, we compare our method with a naive streamable baseline, which trains an independent set of 3D Gaussian primitives for each timestep, referred to as “Naive Per-Timestep 3DGS” in Tab. 5. In this baseline, representation of each timestep is trained with the same conditions as the initial zero-timestep in streaming methods. Although this naive baseline outperforms all off-the-shelf baselines in 4D reconstruction, it fully trains a set of Gaussian primitives from scratch for each scene, leading to a significant training time overhead of approximately 615 seconds per-timestep. In contrast, our DASS effectively models the temporal variations in the 3D space and facilitates a fast convergence between timesteps. DASS demonstrates competitive streaming quality, with only a negligible gap of 0.08 dB compared to this naive baseline, while consuming merely 1.56% of the training time. This significant improvement highlights the efficiency of DASS. We also compare our method with an intuitive full-tuned baseline, which iteratively tunes all parameters of the Gaussian primitives initialized from the previous timestep, termed “Intuitive Full-Tuned 3DGS” in Tab. 5. Specifically, it initializes the on-the-fly training with the same timestep 0 initialization as our DASS and tunes all Gaussian properties for a fixed number of iterations, which yields a similar time consumption to ours. As shown in Tab. 5, a significant performance degradation of 1.16 dB is observed in this baseline. This indicates that the intuitive baseline fails to capture temporal variations between timesteps under such time efficiency, while our DASS effectively manages and converges these variations with lower temporal overhead.
@@
C.3 Time Consumption
The average time consumption for each stage in our DASS framework is provided as follows: Inheritance (0.62 s), Shift (4.02 s), and Densification (1.75 s), with an additional 3.23s spent on fixed expenses, including $\bm{\mathcal{G}}_{0}$ initialization, data loading, and saving. With low training time overhead, our designs efficiently facilitate the transformation of existing elements and retain emerging objects.
C.4 Qualitative Results
We provide additional qualitative results of our method in Fig. 8, Fig. 9, and Fig. 10. Each figure includes the rendered test view visualizations and comparisons with baseline methods. Our DASS presents high-quality reconstruction and fine visual details compared with the baseline methods. We also provide visualizations of the detected dynamics components, which effectively guides the optimization to focus on dynamic objects and complex motions.
Appendix D Analysis on Concurrent Works
We have noticed several concurrent studies on iterative 4D reconstruction, although their focuses and contributions differ from ours. For instance, S4D [17] directly manipulates 3D control points to guide the movements, but it requires full optimization that incurs significant time costs on key frames. V3 [42] aims to facilitate streaming free-viewpoint videos on mobile devices, with a particular emphasis on the reduction of the per-timestep streaming storage overhead. Conversely, our work identifies the per-timestep convergence speed as the critical bottleneck. Moreover, while V3 decomposes motion and appearance attributes for efficient reconstruction, it primarily applies to scenarios with moving objects against vacant background and does not accommodate emerging objects. In addition, SwinGS [28] concentrates on maintaining a consistently stable stream data volume over time and reducing storage overhead through the use of Markov Chain Monte Carlo [22] and a window-based design. Another work [38] considers the bit allocation in a system-level implementation. In contrast, our work prioritizes on-the-fly per-timestep training efficiency as the key bottleneck and focuses on obtaining high-quality streaming while minimizing optimization time.
Appendix E Discussion